# Testing Quantum Gravity with Gravitational Waves from the ringdown of binary Black Holes coalescences : A New Frontier in Fundamental Physics

### v1

@@ 2 Quasi Normal Modes and the Ringdown emission phase
$\frac{d^{2}\Psi_{RW}(r_{*})}{dr_{*}^{2}}+\left[\omega^{2}-V_{RW}(r_{*})\right]\Psi_{RW}(r_{*})=0,$ (2.1)
@@
$V_{RW}(r_{*})=f(r_{*})\left(\frac{l(l+1)}{r_{*}^{2}}-\frac{6GM}{c^{2}\,r_{*}^{3}}\right).$ (2.2)
@@
$V_{Z}(r_{*})=\frac{2f(r_{*})}{r_{*}^{3}(\Lambda r_{*}+3M)^{2}}\left[\Lambda^{2}(\Lambda+1)r_{*}^{3}+3\,\frac{GM}{c^{2}}\Lambda^{2}r_{*}^{2}+9\left(\frac{GM}{c^{2}}\right)^{2}\Lambda r_{*}+9\left(\frac{GM}{c^{2}}\right)^{3}\right],$ (2.3)
@@
For rotating or charged BHs, the Teukolsky equation [26, 27] after the Sasaki–Nakamura transformation [28, 29] reduces to eq. 2.1 with effective potential:
$V_{SN}(r_{*})\simeq\frac{\Delta}{(r_{*}^{2}+a^{2})^{2}}\left[\lambda+\frac{\Delta^{\prime}}{r_{*}^{2}+a^{2}}r_{*}-\frac{3\Delta r_{*}^{2}}{(r_{*}^{2}+a^{2})^{2}}\right],$ (2.4)
@@
$\psi(r_{*})\sim e^{+i\omega r_{*}}\quad r_{*}\rightarrow+\infty,\qquad\psi(r_{*})\sim e^{-i\omega r_{*}}\quad r_{*}\rightarrow-\infty,$ (2.5)
@@
$\omega_{nlm}\simeq\sqrt{V_{0}}\pm i\left(n+\frac{1}{2}\right)\sqrt{2|V_{0}^{\prime\prime}|},$ (2.6)
@@
$\omega_{nlm}\simeq\frac{c^{3}}{GM}\left(\frac{\ln{3}}{8\pi M}+\frac{1}{4M}i\left(n+\frac{1}{2}\right)+O(\sqrt{n})\right),$ (2.7)
@@
$h_{\mu\nu}^{nlm}(t,\,x)=\sum_{n,\,l}A_{\mu\nu}^{\;\;lnm}e^{\left(-\mathfrak{Im}(\omega_{nlm})(t-t_{0})\right)}\cos{\left(\mathfrak{R}(\omega_{nlm})(t-t_{0})\right)},$ (2.8)
@@ 3 Black Holes and Quantum Mechanics
$A=4\,\frac{l_{Pl}^{2}}{k_{B}}\,S=\alpha\,l_{Pl}^{2}\overline{n},$ (3.2)
@@
$\Delta A=\alpha\,l_{Pl}^{2}\,\Delta\overline{n}=\alpha\frac{\hbar G}{c^{3}}\,\Delta\overline{n}$ (3.3)
@@
$\Delta A\simeq\frac{32\pi\,G^{2}}{c^{4}}M\Delta M$ (3.4)
@@
$\Delta E_{\rm Bek}\sim\frac{\hbar c^{3}}{GM},$ (3.5)
@@
$N_{\text{tot}}=\frac{\Delta E_{\rm QNM}}{\Delta E_{\rm Bek}}=10^{-3}\frac{GM^{2}}{\hbar\,c}$ (3.6)
@@
$\frac{\hbar\,\omega_{0}}{\hbar\,\Omega_{\text{QNM}}}=\frac{\omega_{0}}{\Omega_{\text{QNM}}}=\frac{\alpha}{32\,\pi\,k}\,\Delta\overline{n}.$ (3.7)
@@
In the Bekenstein scenario, BH entropy is quantized in levels corresponding to different area eigenstates via eq. 3.1. By contrast, in ST and LQG the entropy is realized as an exponential degeneracy of microstates, with the total number of configurations $\mathcal{N}\simeq e^{S}$ for given entropy value $S$. Therefore, for every macrolevel of entropy a large number of microstates with similar entropy is predicted. Thanks to the correspondence between entropy and mass, this naturally leads to the hypothesis that the energy spectrum is not perfectly sharp but consisting of a large number of degenerate levels with similar energies, concentrated around the energy levels corresponding to the different area eigenstates. The microscopic level spacing inside each band is $\delta E_{micro}\sim\delta E_{M}\,e^{-S}$, where it is supposed to be $\delta E_{M}<\Delta E_{M}$.
@@
$=\int_{0}^{+\infty}\hbar\omega|\omega\rangle\langle\omega|d\omega+\int_{0}^{+\infty}\frac{-i\hbar}{2}\gamma(\omega)|\omega\rangle\langle\omega|d\omega+\int_{0}^{+\infty}\left[g(\omega,\,\omega^{\prime})|\omega\rangle\langle\omega^{\prime}|+\text{h.c.}\right]d\omega\,d\omega^{\prime}$ (3.8)
@@
$|\Psi_{n}\rangle=\int_{\Omega^{n}_{M}-\omega_{n}/2}^{\Omega^{n}_{M}+\omega_{n}/2}c_{n}(\omega)|\omega\rangle d\omega,$ (3.9)
@@
$\int_{\Omega^{n}_{M}-\omega_{n}/2}^{\Omega^{n}_{M}+\omega_{n}/2}|c_{n}(\omega)|^{2}d\omega=1,$ (3.10)
@@
$c_{n}(\omega)\sim\frac{1}{\sqrt{\omega_{n}}}.$ (3.11)
@@
$\mathcal{H}_{P}=P\mathcal{H}P+P\mathcal{Q}\frac{1}{E-Q\mathcal{H}Q}Q\mathcal{H}P=\sum_{n}\alpha_{n}\bigg(\hbar\,\Omega^{n}_{M}+\Theta_{n}(\omega)\bigg)|\Psi_{n}\rangle\langle\Psi_{n}|$ (3.12)
@@
$\Theta_{n}(\omega)=\int\frac{|G_{n}(\omega^{\prime})|^{2}}{\omega-\omega^{\prime}+i0^{+}}d\omega^{\prime}$ (3.13)
@@
$G_{n}(\omega)=\int_{\Omega^{n}_{M}-\omega_{n}/2}^{\Omega^{n}_{M}+\omega_{n}/2}g(\omega,\,\omega^{\prime})c_{n}(\omega^{\prime})d\omega^{\prime}.$ (3.14)
@@
$\Gamma_{n}(\omega)=2\pi\rho(\omega)|G_{n}(\omega)|^{2}.$ (3.15)
@@
$\Gamma_{n}(\omega)\sim g_{0}^{2}\,\hbar\omega_{n},$ (3.16)
@@ 4 Quantum Gravity in the ringdown GW emission
$A=\frac{8\pi M^{2}G^{2}}{c^{4}}\left[1+\sqrt{1-\frac{J^{2}c^{2}}{G^{2}M^{4}}-\frac{Q^{2}}{4\pi\epsilon_{0}\,GM^{2}}}\right]-\frac{GQ^{2}}{\epsilon_{0}\,c^{4}},$ (4.1)
@@
$\Delta A=\frac{16\pi\,G^{2}M}{c^{4}}\left(1+\frac{1}{\sqrt{1-a^{2}}}\right)\Delta M-\frac{8\pi\,Ga}{c^{4}\,\sqrt{1-a^{2}}}\Delta J+\mathcal{O}(Q)$ (4.2)
In a damped oscillator described by the Schrödinger-type eq. 2.1 equation, the ringdown energy depends on the allowed frequencies, which correspond to the poles in the complex plane of the Fourier transform of the Green function of eq. 2.1:
$\left[\frac{d^{2}}{dr_{*}^{2}}+\left(\omega^{2}_{\overline{n}lm}-V_{eff}(r)\right)\right]G(\omega,\,r_{*},\,r_{*}^{\prime})=\delta(r{*}-r_{*}^{\prime}),$ (4.3)
@@
$\Delta M\,c^{2}=\hbar\,|\omega_{\overline{n}lm}|=\hbar\,\sqrt{(\mathfrak{Re}(\omega_{\overline{n}l}))^{2}+(\mathfrak{Im}(\omega_{\overline{n}l}))^{2}}$ (4.4)
@@
Neglecting the charge $Q$, considering $\Delta J=-\hbar m$, substituting $\Delta M$ as obtained respectively in eq. 4.2 in eq. 4.4 and using the quantized area $A$ and $\Delta A$ as defined respectively in eqs. 3.2 and 3.3, the following relation [15] for the allowed ringdown frequencies can be obtained:
$|\omega_{\overline{n}lm}|=\frac{c^{3}}{16\,\pi\,GM}\frac{\overline{n}\alpha\,\sqrt{1-a^{2}}+8\pi\,am}{\sqrt{1-a^{2}}},$ (4.5)
@@
$|\omega_{\overline{n}lm}|=\frac{\alpha\,c^{3}}{32\pi\,GM}\overline{n}.$ (4.6)
@@
$\omega_{\text{QNM}}=\frac{k\,c^{3}}{GM},$ (4.7)
@@
This consideration motivates a simplified illustrative model: a redefinition of the frequencies such that they are spaced proportionally according to eq. 4.5. This model is not derived from a specific quantum gravity theory but rather serves as a heuristic representation of possible quantum effects, based on the hypothesis that the QNM frequencies are quantized. This framework illustrates how a quantum-modified discrete emission spectrum could differ from that predicted by classical QNM theory. Moreover, considering the behavior of their complex parts, Maggiore had already hypothesized that QNM frequencies should be quantized [45]. In fig. 1, we illustrate how the ringdown phase may be modified by redistributing the frequencies so that they become equally spaced. We evaluate the expected ringdown spectrum for a remnant BH with $M=100M_{\odot}$ at a distance of $100\text{Mpc}$. We consider the first 12 modes predicted by the most accurate approach to QNMs, namely the continued fraction solution. The plot is limited to these 12 modes, since for higher orders the damping factor becomes dominant and the corresponding modes are strongly suppressed. In fig. 2, we show the Amplitude Spectral Density (ASD) of the foreseen spectrum for the same ringdown emission cases. For a deterministic signal the ASD is essentially the magnitude of its Fourier transform, normalized per unit frequency, that is the distribution of the signal’s amplitude across frequencies: $ASD(f)=\sqrt{2/T}\,|\tilde{h}(f)|$, where $T$ is the signal duration, and $\tilde{h}(f)$ is the Fourier trasform of the strain as a function of the frequency $f$. Also in this case, the differences induced by the possible quantization scenarios considered are visible.
@@ 5 Quantum Gravity affecting the BH photon sphere
QG models can also be tested in the context of GW emission during the ringdown phase, through the modifications they can introduce on the BH near geometry. This may affect the effective potential related to eq. 2.1 affecting the QNM frequencies or producing echo effects.
@@
$S=\int\sqrt{-g}\left[\frac{c^{3}}{16\pi G}R+l_{Pl}^{2}\,\mathcal{L}_{Pl}+\mathcal{L}_{matt}\right]$ (5.1)
@@
$f(r)=f_{0}(r)+\Delta f(r)=1-\frac{2\,GM}{c^{2}\,r}+\Delta f(r).$ (5.4)
@@
$\Delta f(r)=\frac{2\,GM}{c^{2}\,r}\sum_{n=2}^{\infty}\beta_{n}\left(\frac{l_{Pl}}{r}\right)^{n},$ (5.5)
@@
$V(r)=V_{0}(r)\left(1+\frac{\Delta f(r)}{f_{0}(r)}\right).$ (5.7)
@@
$\frac{dr_{*}}{dr}=\frac{1}{f_{0}(r)+\Delta f(r)}.$ (5.8)
@@
In fig. 3 we report the modifications induced in the Teukolsky potential by quantum gravity perturbations represented in eq. 5.6 form. The plot is obtained for a Shwarzschikd BH remnant of $100\,M_{\odot}$, with $l=2$, considering QG corrections of some percent points of the BH radius, such as foreseen in some BH regularization models [48].
@@
$ds^{2}_{\text{eff}}=e^{2A(r)}\left(-f_{0}(r)\,c^{2}\,dt^{2}+f_{0}^{-1}(r)dr^{2}+r^{2}d\Omega^{2}\right).$ (5.10)
The corresponding Regge–Wheeler eq. 2.2 or Zerilli eq. 2.3 potentials are given by:
$V(r)=f_{\text{eff}}(r)\,\ U(r)$ (5.11)
@@
$V^{\prime}(r)=f^{\prime}_{\text{eff}}(r)U(r)+f_{\text{eff}}U^{\prime}(r)=2A^{\prime}(r)e^{2A(r)}f(r)U(r)+e^{2A(r)}\big[f^{\prime}(r)U(r)+f(r)U^{\prime}(r)\big]=0.$ (5.12)
A similar expression can be found in the case of Teukolsky potential eq. 2.4, hence every BH spacetime can be affected by this modification.
In the cases previously analyzed in eqs. 2.2, 2.3 and 2.4 the potentials featured a single peak, which reflects incoming perturbations while allowing some to transmit into the BH. This process leads to the initial GW emission, known as the primary ringdown signal. In the Schwarzschild case this equation admits a single solution associated with the photon sphere, the inclusion of a non-trivial warp profile introduces additional terms proportional to $A^{\prime}(r)$, which can generate further critical points. In some warp scenarios, therefore, the effective potential can develop a second maximum depending on the contribution of the warp factor $A(r)$. The main peak corresponds to the BH photon sphere and is located outside the EH, whereas the secondary peak is located between the principal maximum and the EH, modifying in an effective way the light ring of the BH. Similar outcomes are found in the contexts of LQG and ST, where quantum modifications to BH EHs are predicted to generate GW echoes [57, 58, 59], even if very weak and not easily detectable. In some models of compact objects the presence of two maxima separated by a minimum forms a cavity that can trap GWs for long periods [54]. This configuration can give rise to long-lived modes, and in certain cases, nonlinear interactions between these modes may cause the trapped energy to grow, potentially destabilizing or even disrupting horizonless objects [55]. The effect is often associated with the presence of a stable light ring, which can confine radiation and enhance the probability of nonlinear behavior. However, as discussed in [56], stable light rings in horizonless spacetimes are typically linked to nonlinear instabilities, whereas BHs —characterized by an absorbing EH— appear to remain dynamically stable configurations. All these models can therefore be tested by searching for late-time echoes in the ringdown phase of GW emission.

### v2

# Testing Quantum Gravity with Gravitational Waves from the Ringdown of Binary Black Hole Coalescences: A New Frontier in Fundamental Physics

@@ Abstract
Keywords: gravitational waves, ringdown, black hole, quantum gravity
@@ 2 Quasi Normal Modes and the Ringdown emission phase
$\frac{d^{2}\Psi_{RW}(r_{*})}{dr_{*}^{2}}+\left[\omega^{2}-V_{RW}(r_{*})\right]\Psi_{RW}(r_{*})=0,$ (1)
@@
$V_{RW}(r_{*})=f(r_{*})\left(\frac{l(l+1)}{r_{*}^{2}}-\frac{6GM}{c^{2}\,r_{*}^{3}}\right).$ (2)
@@
$V_{Z}(r_{*})=\frac{2f(r_{*})}{r_{*}^{3}(\Lambda r_{*}+3M)^{2}}\left[\Lambda^{2}(\Lambda+1)r_{*}^{3}+3\,\frac{GM}{c^{2}}\Lambda^{2}r_{*}^{2}+9\left(\frac{GM}{c^{2}}\right)^{2}\Lambda r_{*}+9\left(\frac{GM}{c^{2}}\right)^{3}\right],$ (3)
@@
For rotating or charged BHs, the Teukolsky equation [26, 27] after the Sasaki–Nakamura transformation [28, 29] reduces to eq. 1 with effective potential:
$V_{SN}(r_{*})\simeq\frac{\Delta}{(r_{*}^{2}+a^{2})^{2}}\left[\lambda+\frac{\Delta^{\prime}}{r_{*}^{2}+a^{2}}r_{*}-\frac{3\Delta r_{*}^{2}}{(r_{*}^{2}+a^{2})^{2}}\right],$ (4)
@@
$\psi(r_{*})\sim e^{+i\omega r_{*}}\quad r_{*}\rightarrow+\infty,\qquad\psi(r_{*})\sim e^{-i\omega r_{*}}\quad r_{*}\rightarrow-\infty,$ (5)
@@
$\omega_{nlm}\simeq\sqrt{V_{0}}\pm i\left(n+\frac{1}{2}\right)\sqrt{2|V_{0}^{\prime\prime}|},$ (6)
@@
$\omega_{nlm}\simeq\frac{c^{3}}{GM}\left(\frac{\ln{3}}{8\pi M}+\frac{1}{4M}i\left(n+\frac{1}{2}\right)+O(\sqrt{n})\right),$ (7)
@@
$h_{\mu\nu}^{nlm}(t,\,x)=\sum_{n,\,l}A_{\mu\nu}^{\;\;lnm}e^{\left(-\mathfrak{Im}(\omega_{nlm})(t-t_{0})\right)}\cos{\left(\mathfrak{R}(\omega_{nlm})(t-t_{0})\right)},$ (8)
@@ 3 Black Holes and Quantum Mechanics
$A=4\,\frac{l_{Pl}^{2}}{k_{B}}\,S=\alpha\,l_{Pl}^{2}\overline{n},$ (10)
@@
$\Delta A=\alpha\,l_{Pl}^{2}\,\Delta\overline{n}=\alpha\frac{\hbar G}{c^{3}}\,\Delta\overline{n}$ (11)
@@
$\Delta A\simeq\frac{32\pi\,G^{2}}{c^{4}}M\Delta M$ (12)
@@
$\Delta E_{\rm Bek}\sim\frac{\hbar c^{3}}{GM},$ (13)
@@
$N_{\text{tot}}=\frac{\Delta E_{\rm QNM}}{\Delta E_{\rm Bek}}=10^{-3}\frac{GM^{2}}{\hbar\,c}$ (14)
@@
$\frac{\hbar\,\omega_{0}}{\hbar\,\Omega_{\text{QNM}}}=\frac{\omega_{0}}{\Omega_{\text{QNM}}}=\frac{\alpha}{32\,\pi\,k}\,\Delta\overline{n}.$ (15)
@@
In the Bekenstein scenario, BH entropy is quantized in levels corresponding to different area eigenstates via eq. 9. By contrast, in ST and LQG the entropy is realized as an exponential degeneracy of microstates, with the total number of configurations $\mathcal{N}\simeq e^{S}$ for given entropy value $S$. Therefore, for every macrolevel of entropy a large number of microstates with similar entropy is predicted. Thanks to the correspondence between entropy and mass, this naturally leads to the hypothesis that the energy spectrum is not perfectly sharp but consisting of a large number of degenerate levels with similar energies, concentrated around the energy levels corresponding to the different area eigenstates. The microscopic level spacing inside each band is $\delta E_{micro}\sim\delta E_{M}\,e^{-S}$, where it is supposed to be $\delta E_{M}<\Delta E_{M}$.
@@
$=\int_{0}^{+\infty}\hbar\omega|\omega\rangle\langle\omega|d\omega+\int_{0}^{+\infty}\frac{-i\hbar}{2}\gamma(\omega)|\omega\rangle\langle\omega|d\omega+\int_{0}^{+\infty}\left[g(\omega,\,\omega^{\prime})|\omega\rangle\langle\omega^{\prime}|+\text{h.c.}\right]d\omega\,d\omega^{\prime}$ (16)
@@
$|\Psi_{n}\rangle=\int_{\Omega^{n}_{M}-\omega_{n}/2}^{\Omega^{n}_{M}+\omega_{n}/2}c_{n}(\omega)|\omega\rangle d\omega,$ (17)
@@
$\int_{\Omega^{n}_{M}-\omega_{n}/2}^{\Omega^{n}_{M}+\omega_{n}/2}|c_{n}(\omega)|^{2}d\omega=1,$ (18)
@@
$c_{n}(\omega)\sim\frac{1}{\sqrt{\omega_{n}}}.$ (19)
@@
$\mathcal{H}_{P}=P\mathcal{H}P+P\mathcal{Q}\frac{1}{E-Q\mathcal{H}Q}Q\mathcal{H}P=\sum_{n}\alpha_{n}\bigg(\hbar\,\Omega^{n}_{M}+\Theta_{n}(\omega)\bigg)|\Psi_{n}\rangle\langle\Psi_{n}|$ (20)
@@
$\Theta_{n}(\omega)=\int\frac{|G_{n}(\omega^{\prime})|^{2}}{\omega-\omega^{\prime}+i0^{+}}d\omega^{\prime}$ (21)
@@
$G_{n}(\omega)=\int_{\Omega^{n}_{M}-\omega_{n}/2}^{\Omega^{n}_{M}+\omega_{n}/2}g(\omega,\,\omega^{\prime})c_{n}(\omega^{\prime})d\omega^{\prime}.$ (22)
@@
$\Gamma_{n}(\omega)=2\pi\rho(\omega)|G_{n}(\omega)|^{2}.$ (23)
@@
$\Gamma_{n}(\omega)\sim g_{0}^{2}\,\hbar\omega_{n},$ (24)
@@ 4 Quantum Gravity in the ringdown GW emission
$A=\frac{8\pi M^{2}G^{2}}{c^{4}}\left[1+\sqrt{1-\frac{J^{2}c^{2}}{G^{2}M^{4}}-\frac{Q^{2}}{4\pi\epsilon_{0}\,GM^{2}}}\right]-\frac{GQ^{2}}{\epsilon_{0}\,c^{4}},$ (26)
@@
$\Delta A=\frac{16\pi\,G^{2}M}{c^{4}}\left(1+\frac{1}{\sqrt{1-a^{2}}}\right)\Delta M-\frac{8\pi\,Ga}{c^{4}\,\sqrt{1-a^{2}}}\Delta J+\mathcal{O}(Q)$ (27)
In a damped oscillator described by the Schrödinger-type eq. 1 equation, the ringdown energy depends on the allowed frequencies, which correspond to the poles in the complex plane of the Fourier transform of the Green function of eq. 1:
$\left[\frac{d^{2}}{dr_{*}^{2}}+\left(\omega^{2}_{\overline{n}lm}-V_{eff}(r)\right)\right]G(\omega,\,r_{*},\,r_{*}^{\prime})=\delta(r{*}-r_{*}^{\prime}),$ (28)
@@
$\Delta M\,c^{2}=\hbar\,|\omega_{\overline{n}lm}|=\hbar\,\sqrt{(\mathfrak{Re}(\omega_{\overline{n}l}))^{2}+(\mathfrak{Im}(\omega_{\overline{n}l}))^{2}}$ (29)
@@
Neglecting the charge $Q$, considering $\Delta J=-\hbar m$, substituting $\Delta M$ as obtained respectively in eq. 27 in eq. 29 and using the quantized area $A$ and $\Delta A$ as defined respectively in eqs. 10 and 11, the following relation [15] for the allowed ringdown frequencies can be obtained:
$|\omega_{\overline{n}lm}|=\frac{c^{3}}{16\,\pi\,GM}\frac{\overline{n}\alpha\,\sqrt{1-a^{2}}+8\pi\,am}{\sqrt{1-a^{2}}},$ (30)
@@
$|\omega_{\overline{n}lm}|=\frac{\alpha\,c^{3}}{32\pi\,GM}\overline{n}.$ (31)
@@
$\omega_{\text{QNM}}=\frac{k\,c^{3}}{GM},$ (32)
@@
This consideration motivates a simplified illustrative model: a redefinition of the frequencies such that they are spaced proportionally according to eq. 30. This model is not derived from a specific quantum gravity theory but rather serves as a heuristic representation of possible quantum effects, based on the hypothesis that the QNM frequencies are quantized. This framework illustrates how a quantum-modified discrete emission spectrum could differ from that predicted by classical QNM theory. Moreover, considering the behavior of their complex parts, Maggiore had already hypothesized that QNM frequencies should be quantized [45]. In fig. 1, we illustrate how the ringdown phase may be modified by redistributing the frequencies so that they become equally spaced. We evaluate the expected ringdown spectrum for a remnant BH with $M=100M_{\odot}$ at a distance of $100\text{Mpc}$. We consider the first 12 modes predicted by the most accurate approach to QNMs, namely the continued fraction solution. The plot is limited to these 12 modes, since for higher orders the damping factor becomes dominant and the corresponding modes are strongly suppressed. In fig. 2, we show the Amplitude Spectral Density (ASD) of the foreseen spectrum for the same ringdown emission cases. For a deterministic signal the ASD is essentially the magnitude of its Fourier transform, normalized per unit frequency, that is the distribution of the signal’s amplitude across frequencies: $ASD(f)=\sqrt{2/T}\,|\tilde{h}(f)|$, where $T$ is the signal duration, and $\tilde{h}(f)$ is the Fourier trasform of the strain as a function of the frequency $f$. Also in this case, the differences induced by the possible quantization scenarios considered are visible.
@@ 5 Quantum Gravity affecting the BH photon sphere
QG models can also be tested in the context of GW emission during the ringdown phase, through the modifications they can introduce on the BH near geometry. This may affect the effective potential related to eq. 1 affecting the QNM frequencies or producing echo effects.
@@
$S=\int\sqrt{-g}\left[\frac{c^{3}}{16\pi G}R+l_{Pl}^{2}\,\mathcal{L}_{Pl}+\mathcal{L}_{matt}\right]$ (33)
@@
$f(r)=f_{0}(r)+\Delta f(r)=1-\frac{2\,GM}{c^{2}\,r}+\Delta f(r).$ (36)
@@
$\Delta f(r)=\frac{2\,GM}{c^{2}\,r}\sum_{n=2}^{\infty}\beta_{n}\left(\frac{l_{Pl}}{r}\right)^{n},$ (37)
@@
$V(r)=V_{0}(r)\left(1+\frac{\Delta f(r)}{f_{0}(r)}\right).$ (39)
@@
$\frac{dr_{*}}{dr}=\frac{1}{f_{0}(r)+\Delta f(r)}.$ (40)
@@
In fig. 3 we report the modifications induced in the Teukolsky potential by quantum gravity perturbations represented in eq. 38 form. The plot is obtained for a Shwarzschikd BH remnant of $100\,M_{\odot}$, with $l=2$, considering QG corrections of some percent points of the BH radius, such as foreseen in some BH regularization models [48].
@@
$ds^{2}_{\text{eff}}=e^{2A(r)}\left(-f_{0}(r)\,c^{2}\,dt^{2}+f_{0}^{-1}(r)dr^{2}+r^{2}d\Omega^{2}\right).$ (42)
The corresponding Regge–Wheeler eq. 2 or Zerilli eq. 3 potentials are given by:
$V(r)=f_{\text{eff}}(r)\,\ U(r)$ (43)
@@
$V^{\prime}(r)=f^{\prime}_{\text{eff}}(r)U(r)+f_{\text{eff}}U^{\prime}(r)=2A^{\prime}(r)e^{2A(r)}f(r)U(r)+e^{2A(r)}\big[f^{\prime}(r)U(r)+f(r)U^{\prime}(r)\big]=0.$ (44)
A similar expression can be found in the case of Teukolsky potential eq. 4, hence every BH spacetime can be affected by this modification.
In the cases previously analyzed in eqs. 2, 3 and 4 the potentials featured a single peak, which reflects incoming perturbations while allowing some to transmit into the BH. This process leads to the initial GW emission, known as the primary ringdown signal. In the Schwarzschild case this equation admits a single solution associated with the photon sphere, the inclusion of a non-trivial warp profile introduces additional terms proportional to $A^{\prime}(r)$, which can generate further critical points. In some warp scenarios, therefore, the effective potential can develop a second maximum depending on the contribution of the warp factor $A(r)$. The main peak corresponds to the BH photon sphere and is located outside the EH, whereas the secondary peak is located between the principal maximum and the EH, modifying in an effective way the light ring of the BH. Similar outcomes are found in the contexts of LQG and ST, where quantum modifications to BH EHs are predicted to generate GW echoes [57, 58, 59], even if very weak and not easily detectable. In some models of compact objects the presence of two maxima separated by a minimum forms a cavity that can trap GWs for long periods [54]. This configuration can give rise to long-lived modes, and in certain cases, nonlinear interactions between these modes may cause the trapped energy to grow, potentially destabilizing or even disrupting horizonless objects [55]. The effect is often associated with the presence of a stable light ring, which can confine radiation and enhance the probability of nonlinear behavior. However, as discussed in [56], stable light rings in horizonless spacetimes are typically linked to nonlinear instabilities, whereas BHs —characterized by an absorbing EH— appear to remain dynamically stable configurations. All these models can therefore be tested by searching for late-time echoes in the ringdown phase of GW emission.